Operation Summary for June 30, 2010 - July 2, 2010
Wednesday June 30
The day shift began with Operations monitoring the Tevatron (TeV) store 7939 (124.57E30) and the Recycler stash (208.3E10), with the Antiproton Source (Pbar) stacking (11.1E10), with the Linac, Booster and the Main Injector (MI) operating normally, with the NuMI and MiniBooNE experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments using the colliding beams.
During the evening shift, Operations established store 7940 at 4:04 PM with an Average Initial Luminosity (AIL) of 325.7E30.
Thursday July 1
The midnight shift began with Operations monitoring store 7940 (98.13E30) and the stash (227.8E10), with Pbar stacking (126.4E10), and with NuMI and MiniBooNE taking beam.
During the midnight shift, Operations established beam to the Meson FTBF T-977 experiment at 4:20 AM. Operations terminated store 7940 at 5:52 AM. Operations put the TeV into shot setup and established store 7941 at 7:03 AM with an AIL of 257.4E30.
During the day shift, Preaccelerator personnel switched to the I- Source at 10:18 AM. Operations terminated store 7941 at 10:35 AM. TeV personnel found that the sector C49 separator polarity switch was the cause of the luminosity jumps; they began repairs. Operators prepared CDF and D0 for accesses.
During the evening shift, Operations put the TeV into shot setup at 4:06 PM and established store 7945 at 5:17 PM with an AIL of 225.6E30.
Friday July 2
The midnight shift began with Operations monitoring store 7945 (83.97E30) and the stash (170.0E10), with Pbar stacking (16.8E10), and with NuMI and MiniBooNE taking beam.
During the midnight shift, Operations established beam to T-977 at 04:09 AM.
The Plans for Friday and the Weekend
The plans for today and the weekend are to stack, stash, and store.
Comments and Suggestions
|last modified 07/02/2010 email Fermilab|
|Security, Privacy, Legal|